The instruction manual on the industrial oven suggests your oven temperature will increase by 45 degrees Fahrenheit per minute. When the over is initially turned on, the temperature is 70 degrees Fahrenheit. What will the temperature of the oven be after 7 minutes?
Found 2 solutions by mangopeeler07, ikleyn:
Answer by mangopeeler07(462)

70+7(45)------it starts at 70 and increases by 45 per minute for 7 minutes

Distribute
70+315

Add
70+315=385

Temp. after 7 minutes:
385 degrees

Answer by ikleyn(52794)

.
The general formula for the temperature, according to the given condition, is


    T = 70 + 45*t   degrees of Fahrenheit.


where "t " is the time in minutes.


Atter 7 minutes  T = 70 + 45*7  degrees of Fahrenheit.

Solved.